# Toward Storage-Aware Learning with Compressed Data An Empirical Exploratory Study on JPEG

### v1

@@ 4 Discussion and Future Works
Synergy with Other On-Device Learning Paradigms. Our storage-aware framework is inherently suited for systems with continuous data streams, creating a natural synergy with paradigms like Federated Learning (FL), Continual Learning (CL), and Active Learning (AL) [8, 9, 14]. Our approach can enhance these methods by providing a more sophisticated data retention policy, for example, by enabling efficient knowledge preservation in CL or managing client data under heterogeneous storage in FL. Integrating our quality-quantity balancing strategy into these frameworks presents a significant opportunity to develop more robust and efficient on-device intelligence.
@@
Addressing the Training Overhead. While our primary focus is the storage budget, we acknowledge that retaining compressed data, rather than discarding it, may increase computational overhead during training. We posit this is a manageable trade-off, as on-device training often occurs during device idle periods [15]. Moreover, this overhead could be actively mitigated. Future work could explore hybrid policies that combine data compression with selective dropping or employ data consolidation techniques to form more efficient training batches, thereby optimizing the balance between storage efficiency and computational cost [16].

### v2

@@ 4 Discussions, Limitations and Future Work
Synergy with Other On-Device Learning Paradigms. Our storage-aware approach is inherently suited for systems with continuous data streams, creating a natural synergy with paradigms like Federated Learning (FL), Continual Learning (CL), and Active Learning (AL) [wang2022memory, lee2024detrigger]. Specifically, our approach enhances these methods by providing a more sophisticated data retention policy, for example, by enabling efficient knowledge preservation in CL or managing client data under heterogeneous storage in FL. Integrating quality-quantity balancing into these frameworks presents a significant opportunity to develop more robust and efficient on-device intelligence.
@@
Addressing the Training Overhead. While our primary focus is the storage budget, we acknowledge that retaining compressed data, rather than discarding it, may increase computational overhead during training. We posit this is a manageable trade-off, as on-device training often occurs during device idle periods. Moreover, this overhead could be actively mitigated. Future work could explore hybrid policies that combine data compression with selective dropping or employ data consolidation techniques to form more efficient training batches, thereby optimizing the balance between storage efficiency and computational cost [zhao2020dataset].